### v2

# Cost-Free Neutrality for the River Method

Abstract
Recently, the River Method was introduced as novel refinement of the Split Cycle voting rule. The decision-making process of River is closely related to the well established Ranked Pairs Method. Both methods consider a margin graph computed from the voters’ preferences and eliminate majority cycles in that graph to choose a winner. As ties can occur in the margin graph, a tiebreaker is required along with the preferences. While such a tiebreaker makes the computation efficient, it compromises the fundamental property of neutrality: the voting rule should not favor alternatives in advance.
One way to reintroduce neutrality is to use Parallel-Universe Tiebreaking (PUT), where each alternative is a winner if it wins according to any possible tiebreaker. Unfortunately, computing the winners selected by Ranked Pairs with PUT is $\mathtt{NP}$-complete. Given the similarity of River to Ranked Pairs, one might expect River to suffer from the same complexity.
Surprisingly, we show the opposite: We present a polynomial-time algorithm for computing River winners with PUT, highlighting significant structural advantages of River over Ranked Pairs. Our Fused-Universe (FUN) algorithm simulates River for every possible tiebreaking in one pass. From the resulting FUN diagram one can then directly read off both the set of winners and, for each winner, a certificate that explains how this alternative dominates the others.
1 Introduction and Related Work
A common interest in theoretical computer science, economics, and political science is the design and analysis of social choice functions-mechanisms that aggregate individual preferences to make a collective decision. Clearly, such decision-making is an integral part of democratic processes. Beyond political contexts, social choice functions have become increasingly relevant in artificial intelligence, particularly for aligning AI systems with diverse human preferences (Fürnkranz and Hüllermeier 2003; Köpf et al. 2024). As language models and other AI systems interact with feedback from a wide range of users, aggregating that feedback ‘socially acceptably’ becomes a key challenge. Recent work argues that social choice theory should guide AI alignment in addressing this diversity (Conitzer et al. 2024).
There is extensive research on properties of social choice functions, their trade-offs, and the often high complexity of computing them (see (Brandt et al. 2016) for an overview).
A core task of computational social choice is to design rules that satisfy reasonable fairness criteria while also being computationally tractable. In this work, we show how to efficiently compute the winners of the recently suggested River Method (Döring et al. 2025).
1.1 Margin-Based Social Choice Functions
River is part of a family of margin-based social choice functions – methods that decide an election based on pairwise comparisons between alternatives and the margin of victory in each such comparison. A common approach to social choice is to select an alternative that wins all pairwise comparisons; the so-called Condorcet winner. However, there are cases where there is no such alternative. Margin-based rules resolve this by not just considering who wins in a pairwise comparison, but also by how much.
The margin of $x$ over $y$ is the difference between the number of voters who prefer $x$ over $y$ and those who prefer $y$ over $x$. This information is represented as a margin graph – a complete, weighted, antisymmetric, directed graph over the set of alternatives, where each edge $(x,y)$ is weighted by the margin between $x$ and $y$ (Brandt et al. 2016). If there is no Condorcet winner, the margin graph contains cycles of pairwise victories, such as alternatives $x,y,z$ where $x$ wins against $y$, which wins against $z$, which wins against $x$.
One way of resolving such cycles is to focus on the notion of immunity to majority complaints: an alternative $x$ is immune if for every alternative $z$ that beats $x$ in direct comparison, there exists a path of at least as strong pairwise victories leading from $x$ to $z$. This path does not only serve to define a winner, but also to defend the choice of $x$ against claims such as “a majority prefers $z$ over $x$, thus $x$ should not win”.
In every election there exists at least one immune alternative (Holliday and Pacuit 2023a). Among the numerous social choice functions selecting immune alternatives, we want to highlight five that are particularly notable for their distinct methods of resolving majority cycles:
  • Split Cycle (Holliday and Pacuit 2023a) aims to identify a consistent winner set by deleting the weakest link in every cycle in the margin graph. As Split Cycle selects all immune alternatives as winners, the following methods can be seen as refinements.
  • Ranked Pairs (Tideman 1987) incrementally locks in pairwise victories, prioritizing those with the largest margins that do not create cycles.
  • Stable Voting (Holliday and Pacuit 2023b) recursively refines the winner set, ensuring stability under the deletion of alternatives.
  • Beat Path (Schulze 2011, 2025) selects winners based on the strongest paths through the defeat graph.
  • River (Döring et al. 2025) builds on Ranked Pairs by providing a simpler approach to refine the majority cycle elimination process.
For an overview of the definitions and properties of these methods, we refer to (Holliday and Pacuit 2023a; Döring et al. 2025).
Among these margin-based choice functions, Ranked Pairs, Stable Voting, and River share a key feature: they process the edges of the margin graph in decreasing order of the margin. When multiple edges have the same margin, these methods require a tiebreaking rule (a total descending linear ordering of the margin edges) as additional input.
Unfortunately, such a tiebreaker usually compromises the property of neutrality, which requires that no alternative is favored a priori. See (Brill and Fischer 2012; Freeman et al. 2015; Wang et al. 2019; Zavist and Tideman 1989) for a discussion of possible tiebreaking rules and their implications. To reintroduce neutrality, one can compute the winning set using the Parallel-Universe Tiebreaking (PUT), which considers all possible tiebreaking orders and returns the union of winners. Unfortunately, deciding whether an alternative is a winner under PUT for Ranked Pairs is known to be $\mathtt{NP}$-complete (Brill and Fischer 2012).
1.2 Our Contribution
We study the computational complexity of applying Parallel-Universe Tiebreaking (PUT) to River. Although River closely resembles Ranked Pairs, we show that the winners under PUT can be computed in polynomial time.
For this, we exploit the more restricted rule set of River: like Ranked Pairs, it processes the edges of the margin graph in order of decreasing margin and discards those that would create a cycle; River additionally enforces that each alternative has at most one incoming edge. As a result, River always produces a directed tree rooted at the winner, whereas Ranked Pairs only guarantees to return an acyclic subgraph.
This difference enables us to simultaneously simulate River for all possible tiebreakers in one pass. Our Fused-Universe ($\mathsf{FUN}$) algorithm to determine all PUT winners for River can be summarized as follows:
  1. 1.
    Compute the Fused-Universe diagram: a subgraph of the margin graph, which includes all margin edges that appear in the River diagram of at least one universe.
  2. 2.
    During that same process, track and update for each alternative whether it remains undefeated in all universes, is defeated in some, or is defeated in all universes.
  3. 3.
    Derive the set of winners from these vertex states.
After introducing some notation and formally defining River in Section 2, we describe our algorithm in detail in Section 3, where we also explain how it can be computed in polynomial time. In Section 4 we prove that the algorithm correctly determines the set of $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winners. We first show that every alternative not selected by our algorithm is defeated in every universe. For the other direction, we construct, for each selected winner, a tiebreaker under which it wins in River. This tiebreaker not only proves correctness but also serves as a constructive certificate that the alternative is immune. We conclude in Section 6 with open questions and suggestions for future work on River and Parallel-Universe Tiebreaking.
2 Preliminaries
We consider $n\geq 1$ voters expressing preferences over a set $A$ of alternatives. The preferences of each voter $i$ are represented by a linear order $\succ_{i}$} over $A$, and the (preference) profile $\mathcal{P}={(\succ_{1},\dots,\succ_{n})}$ collects these preferences for all voters. Let $n]=\{1,\dots,n\}$. The margin of alternative $x$ over $y$ according to $\mathcal{P}$ is $m_{\mathcal{P}}(x,y)=\left\lvert\{\,i\in[n]\mid x\succ_{i}y\,\}\right\rvert-\left\lvert\{\,i\in[n]\mid y\succ_{i}x\,\}\right\rvert$.
The margin graph $\mathcal{M}(\mathcal{P})$ of $\mathcal{P}$ is a weighted directed graph with vertex set $A$ and edge set $\{\,(x,y)\in A\times A\mid m_{\mathcal{P}}(x,y)\geq 0\,\}$, where each edge $(x,y)$ is assigned weight $m_{\mathcal{P}}(x,y)\geq 0$. In any graph with vertex set $A$, we say that vertex $y$ is dominated by the edge $(x,y)$ or by vertex $x$, if $y$ has an incoming edge from $x$. We omit the preference profile $\mathcal{P}$ when it is clear from the context; for example, we write $m(x,y)$ for a margin function and $\mathcal{M}$ for a margin graph.
A (majority) path from $x$ to $y$ in $\mathcal{M}$ is a sequence $P_{xy}=(x=p_{1},\ldots,p_{\ell}=y)$ of distinct alternatives such that $m(p_{i},p_{i+1})>0$ for all $i\in[\ell-1]$. The strength of such a path is the value of its lowest margin, i. e., $\operatorname{strength}(P_{xy})=\min\{\,m(p_{i},p_{i+1})\mid i\in[\ell-1]\,\}$. Analogously, majority cycles in $\mathcal{M}$ are closed paths ($x_{\ell}=x_{1}$) such that $m(x_{i},x_{i+1})>0$ for all consecutive pairs, and their strength is defined as the lowest margin on the cycle.
Winners and Ties
A Condorcet winner $x$ is an alternative which defeats all other alternatives in a pairwise comparison, i. e., $m(x,y)>0$ for all $y\in A\setminus\{x\}$. Such an alternative does not always exist. In the absence of a Condorcet winner, every alternative faces at least one majority defeat. Some alternatives can be justified as winners despite these defeats by the existence of equally strong majority paths leading from the alternative to each alternative defeating it. This notion is called immunity (Holliday and Pacuit 2023a): An alternative $x$ is immune (against majority complaints) if, for every alternative $y$ with $m(y,x)>0$, there is a majority path $P_{xy}$ with $\operatorname{strength}(P_{xy})\geq m(y,x)$.
[…]
The River Method
River (Döring et al. 2025) is a social choice function which eliminates majority cycles in the margin graph by greedily constructing a subtree of $\mathcal{M}$ based on a fixed tiebreaker. Formally, given a preference profile $\mathcal{P}$ and a tiebreaker $\tau\in\mathbb{L}$, let $\mathcal{M}=(A,E)$ denote the corresponding margin graph. River constructs the River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$, and selects a unique winner $\mathsf{RV}(\mathcal{P},\tau)$ as follows:
  1. 1.
    Initialize $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$ as $(A,\emptyset)$.
  2. 2.
    Iterate over the edges in order of $\tau$ and add $(x,y)$ to the diagram unless it is rejected by one of two conditions:
    branching condition:
    $y$ already has an incoming edge;
    cycle condition:
    adding the edge would create a cycle.
  3. 3.
    Return the unique alternative without incoming edges in $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$ as winner $\mathsf{RV}(\mathcal{P},\tau)$.
Parallel-Universe Tiebreaking
Instead of fixing one tiebreaker, Parallel-Universe Tiebreaking (PUT) selects every winner according to any possible tiebreaker. Formally,
$\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})=\bigcup_{\tau\in\mathbb{L}}\mathsf{RV}(\mathcal{P},\tau).$
3 Fusing Parallel Universes
The $\mathsf{FUN}$ algorithm efficiently simulates the River procedure across all universes simultaneously. To do so, it exploits River’s unique branching condition: every alternative has at most one incoming edge in any River diagram. This allows us to reason about which edges might be added or rejected across all possible universes.
Consider the margin graph in Figure 1. The edges $(x,y)$, $(z,y)$, and $(d,y)$, with margin $10$, are processed first by River. Depending on the tiebreaker, one is added, while the other is rejected by the branching condition. As a result, any edge to $y$ with margin lower than 10 is rejected by the branching condition regardless of the tiebreaker. This allows us to reason about potential winners: once all margin-10 edges are processed, it is already certain that $y$ cannot be a River winner in any universe. Indeed, one of $(x,y)$, $(z,y)$, or $(d,y)$ will appear in every River diagram, thereby dominating $y$. Consequently, $y$ cannot be an $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winner.
short edge state interpretation implication
Fix fix added in every universe in every universe, $y$ is dominated by $(x,y)$
BC branching choice added in some but not all universes; there is $(z,y)$ with $m(z,y)=k$ which could be added instead in every universe, $y$ is dominated by an edge with margin $k$
$C$ cycle choice may be added in some but not all universes; there exists an $y$-$x$-path which could be added instead there is a universe where $y$ is not dominated by $(x,y)$
CBC cycle branching choice added in some but not all universes; there exists $(z,y)$ with $m(z,y)>k$ which is CC in every universe, $y$ is dominated by an edge with margin at least $k$
Table 1: Overview of the possible edge states assigned to an edge $(x,y)$ with margin $k=m(x,y)$.
The $\mathsf{FUN}$ algorithm. Our algorithm to compute the $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winners has two main steps.
In the first step, it computes the Fused-Universe ($\mathsf{FUN}$) diagram. The algorithm takes as input the margin graph and processes its edges in order of decreasing margin; edges with the same margin are considered in arbitrary order. For each edge $(x,y)$ with margin $k$ the algorithm checks whether the edge is clearly rejected from every possible universe because of River’s branching or cycle condition. If the algorithm has decided that $(x,y)$ cannot be clearly rejected from every universe, it adds the edge to the $\mathsf{FUN}$ diagram and proceeds to compute the conditions under which this edge may appear in some universe. These conditions are represented by four possible edge states: fix, branching choice, cycle choice, or cycle-branching choice. Each alternative – represented in the diagram as a vertex – also receives a vertex state: not dominated, fixedly dominated, or cycle dominated.
In the second step, the $\mathsf{FUN}$ algorithm returns the set of winning alternatives, which is the set of vertices which are not dominated or cycle dominated in the $\mathsf{FUN}$ diagram.
We introduce the edge and vertex states in Section 3.1, followed by a detailed description of their computation in Section 3.2. Afterwards, we explain in Section 3.3 how the algorithm can be implemented to run in polynomial time.
3.1 Edge and Vertex States
Edge States
If run with a fixed tiebreaker, River either rejects or adds an edge to the River diagram. Our algorithm is more nuanced about the edge states, as it considers simultaneously whether edges are rejected or added in any possible universe. An edge in the $\mathsf{FUN}$ diagram receives one of four edge states– fix, branching choice, cycle choice, and cycle branching choice. Their interpretations, implications, and abbreviations are summarized in Table 1 and illustrated in Figure 1. Let $(x,y)$ with margin $k$ be the current edge.
A fix edge appears in every universe. An edge is marked Fix if we are certain that no other incoming edge to $y$ of margin $\geq k$ and no path from $y$ to $x$ has strength $\geq k$. A simple example for a Fix edge is a unique edge with maximum margin: since that edge is considered as the very first edge in every universe, neither River’s branching nor cycle condition reject it and the edge is added to every River diagram.
A branching choice edge may appear in some universe but cannot appear in all. That is because there exists at least one other incoming “partner” edge $(z,y)$ with margin $k$ which also could be added. Which of the two edges is added to the River diagram depends on the order in the tiebreaker $\tau$: if $(x,y)$ is before $(z,y)$ in $\tau$, then $(x,y)$ is added and $(z,y)$ must be rejected by the branching condition. Vice versa, if $(z,y)$ is before $(x,y)$ in $\tau$, then $(z,y)$ is added and $(x,y)$ rejected. Both edges are marked as a branching choice BC.
A cycle choice edge may appear in some but not all universes. Assume there is a path from $d$ to $c$ containing at least one BC edge $(d,y)$, and let $(z,y)$ be the partner BC that is not on the path. If the tiebreaker puts $(z,y)$ first, $(d,y)$ is rejected by the branching condition. Thus, the addition of $(c,d)$ does not close a cycle and the edge is added. If $(d,y)$ is put first instead, the addition of $(c,d)$ would close a cycle, and it would be rejected by the cycle condition. Another simple example would be a directed cycle $(a,b,c)$ containing only edges of margin $k$. Depending on the order in the tiebreaker, two of the three edges will be added to the River diagram while the third is rejected by the cycle condition.
A cycle branching choice edge is a branching choice which is dependent on the cycle choice of another edge. Assume $(c,d)$ is a cycle choice edge because there is a path from $d$ to $c$ that includes a branching choice edge $(d,y)$ with partner $(z,y)$ (not on the path). In that case, the addition of $(x,d)$ depends on the choices made on the cycle and might lead to a rejection by the branching condition: if $(z,y)$ is processed first and added, then $(d,y)$ is rejected, $(c,d)$ does not close a cycle and is added; consequently, $(x,d)$ must be rejected by the branching condition. If $(d,y)$ is processed and added first instead, then $(c,d)$ is rejected by the cycle condition and $(x,d)$ is added.
Vertex States
vertex state property in $\mathcal{M}^{\mathsf{FUN}}$}
not dominated no incoming edge
fixedly dominated either exactly one incoming edge with state Fix, or at least one incoming edge with state BC or CBC
cycle dominated at least one incoming edge and all incoming edges have state CC
Table 2: Overview of the possible vertex states.
Figure 1: Election with margin graph $\mathcal{M}$, $\mathsf{FUN}$ diagram $\mathcal{M}^{\mathsf{FUN}}$} with edge/vertex states and $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winner $x$.
If run with a fixed tiebreaker, River selects a winner after all edges have been processed – the unique alternative with no incoming edge in the resulting River diagram. Our algorithm generalizes this by assigning vertex states during execution, capturing structural information about which alternatives can or cannot be winners across all universes.
These vertex states reflect what is known about a vertex $y$’s incoming edges in $\mathcal{M}^{\mathsf{FUN}}$}. Edges labeled Fix, BC, or CBC with margin $k$ indicate that $y$ is dominated by an edge of margin $\geq k$ in every universe. In contrast, CC edges do not guarantee such domination, as they may be rejected depending on the cycle structure and tiebreaker. This distinction allows us to assign each vertex one of three possible states, as summarized in Table 2 and illustrated in Figure 1.
3.2 Computing the Fused-Universe Diagram
Let us now go into the details on how to compute the introduced edge and vertex states for the $\mathsf{FUN}$ diagram. Given a margin graph $\mathcal{M}$, the algorithm processes the edges in descending order of their margin, breaking ties arbitrarily. Algorithm 1 shows how each edge $(x,y)$ with margin $m(x,y)=k$ is then processed in two phases:
Input: margin graph $\mathcal{M(\mathcal{P})}=(A,E)$ of a preference profile $\mathcal{P}$ with margins $m$
Output: $\mathsf{FUN}$ diagram $\mathcal{M}^{\mathsf{FUN}}$}
1ex
${\mathcal{M}^{\mathsf{FUN}}}\leftarrow(A,\emptyset)$
1exforeach $(x,y)\in E$ in order of decreasing margin do
 $k\leftarrow m(x,y)$
 // branching reject check
 if $y$ is fixedly dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then reject $(x,y)$ and continue
 // cycle reject check
 if there is a path from $y$ to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$} then
      compute the set of vertices $U$ that have a path to $x$ of strength >k>k in $\mathcal{M}^{\mathsf{FUN}}$}
    if for all $u\in U$ there exists a path $P\in{\mathcal{M}^{\mathsf{FUN}}}$ from $y$ to $u$ of strength >k>k, such that for every CC edge $(c,d)\in P$, all paths from $d$ to $c$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength $\geq m(c,d)$ include vertex $y$ then
       reject $(x,y)$ and continue
      end if
   end if
 // $(x,y)$ will be added to $\mathcal{M}^{\mathsf{FUN}}$} – compute tentative state of $(x,y)$
 if $y$ is not dominated then $state(x,y)\leftarrow\textnormal{{Fix}}$
 else if $y$ is cycle dominated and $\exists(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)>k$ then $state(x,y)\leftarrow\textsf{CBC}$
 else if $y$ is cycle dominated and $\forall(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}\colon m(z,y)=k$ then $state(x,y)\leftarrow\textsf{BC}$
 else // $y$ is fixedly dominated by edges of margin $k$
    $state(x,y)\leftarrow\textsf{BC}$
      update incoming Fix edge of $y$ to BC
   end if
 add $(x,y)$ to $\mathcal{M}^{\mathsf{FUN}}$}
 // cycle update check
  compute the set of edges $C\subseteq{\mathcal{M}^{\mathsf{FUN}}}$ that form a cycle with $(x,y)$
 if $C\neq\emptyset$ then
    $state(x,y)\leftarrow\textsf{CC}$
    foreach $e\in C$ with $m(e)=k$ do $state(e)\leftarrow\textsf{CC}$
   end if
end foreach
return $\mathcal{M}^{\mathsf{FUN}}$}
Algorithm 1 Computing the fused universe diagram.
1. Rejection in Every Universe (lines 1 to 1).
The algorithm first decides whether the edge $(x,y)$ must be rejected from $\mathcal{M}^{\mathsf{FUN}}$} in every tiebreaker universe. This decision is based on River’s branching and cycle conditions:
  • Branching rejection: If vertex $y$ is fixedly dominated by some edge with margin >k>k, then $(x,y)$ is rejected by River’s branching condition in every universe.
  • Cycle rejection: If adding $(x,y)$ forms a cycle with a path of strength >k>k in every universe, then in each of these universes the edge is rejected by River’s cycle condition. This is checked by the algorithm in lines 11.
    The condition in line 1 can be computed with a breadth-first search (BFS) started on $y$ that only considers edges with margin >k>k. Similarly, the set $U$ in line 1 can be computed with a reverse BFS started on $x$ that only considers edges with margin >k>k. To check the condition in line 1, we do the following. For every CC edge $(c,d)$ in $\mathcal{M}^{\mathsf{FUN}}$}, run a BFS started on $d$ that only considers edges with margin $\geq m(c,d)$ that are not incident to $y$. Let $E^{\prime}$ be the set of such edges $(c,d)$ where this BFS reaches $c$ – meaning there is a path from $d$ to $c$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength $\geq m(c,d)$ that does not include vertex $y$. Now the condition in line 1 can be tested by comparing $U$ with the vertices reached by a BFS started on $y$ that only considers edges in $E\setminus E^{\prime}$ with margin >k>k.
2. Assigning the Edge State (lines 1 to 1).
If the edge is not rejected in Step 1, the algorithm will add the edge to $\mathcal{M}^{\mathsf{FUN}}$} and determine its state. A first preliminary state is derived from the current state of the target vertex $y$, without yet accounting for possible cycles in $\mathcal{M}^{\mathsf{FUN}}$} – resulting in one of the states Fix, BC, or CBC (lines 11). Later, the algorithm checks whether $(x,y)$ is contained in cycles in $\mathcal{M}^{\mathsf{FUN}}$} to determine whether it (and possibly other edges of margin $k$) must be updated to CC (lines 11).
Note that the state of an edge is tentative until all edges of the same margin are processed. Only then it is clear which cycles this edge is contained in. For this, in lines 11, the algorithm performs the cycle update check: it updates the edges in cycles that are closed by $(x,y)$ in $\mathcal{M}^{\mathsf{FUN}}$}. To compute CC in 1, the algorithm performs two BFS in $\mathcal{M}^{\mathsf{FUN}}$}: A forward BFS from $y$ to find all vertices $Y$ reachable from $y$ and a reverse BFS from $x$ to find all vertices $X$ reaching $x$. Now CC is the set of edges $(u,v)$ with $u$ in $Y$ and $v$ in $X$.
3.3 Computing the $\operatorname{\mathsf{RV\mbox{-}PUT}}$ Winners in Polynomial Time
After computing the Fused-Universe diagram $\mathcal{M}^{\mathsf{FUN}}$}, our algorithm returns the set of vertices that are cycle dominated or not dominated in $\mathcal{M}^{\mathsf{FUN}}$} as $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winners. These two steps run in polynomial time in the size of the margin graph.
Theorem 3.1.
The runtime of the $\mathsf{FUN}$ algorithm executed on a margin graph $\mathcal{M}=(A,E)$ is polynomial in the number of alternatives $|A|$.
Proof.
We assume that $\mathcal{M}$ is given as adjacency lists. Likewise, the algorithm constructs $\mathcal{M}^{\mathsf{FUN}}$} using adjacency lists and stores each edge state along with the respective edge.
Observe that $|E|\in\Theta(|A|^{2})$ and that the following operations can be computed in polynomial time in $|A|$ on both graphs: (1) initializing and adding edges to $\mathcal{M}^{\mathsf{FUN}}$}; (2) sorting and iterating through all edges; (3) iterating over incoming edges of a vertex; (4) determining the state of a vertex; (5) running a breadth first search (BFS); (6) transposing the graph and running a reverse BFS; (7) updating an edge state; (8) iterating over all vertices.
This covers all elemental operations of Algorithm 1 as well as the final pass over all vertices to determine the winner set. In Section 3.2 we explain how the complex operations (namely cycle reject check and cycle update check) reduce to running $\mathcal{O}(|E|)$ breadth first searches. Thus, the $\mathsf{FUN}$ algorithm overall runs in polynomial time in $|A|$. ∎
4 The $\mathsf{FUN}$ Algorithm Returns Exactly the $\operatorname{\mathsf{RV\mbox{-}PUT}}$ Winners
In the following two sections, we prove that one can correctly identify the $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winners using the $\mathsf{FUN}$ diagram. Formally, we show the following theorem.
Theorem 4.1.
For every preference profile $\mathcal{P}$ and every alternative $a\in A$: $a\in\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})\Leftrightarrow a$ is cycle dominated or not dominated in ${\mathcal{M}^{\mathsf{FUN}}}(\mathcal{P})$.
4.1 Forward Direction: Fix, Branching Choice and Rejected Edges
In this section, we show the forward direction of Theorem 4.1 which states that every $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winner $a$ is identified by our $\mathsf{FUN}$ algorithm. To that end, we first establish a structural lemma relating the state of an edge in the $\mathsf{FUN}$ diagram to its presence or absence in the River diagram under any tiebreaker. We write $\mathcal{M}^{\mathsf{RV}}_{\tau}$ as a shortcut for the River diagram $\mathsf{RV}(\mathcal{P},\tau)$ of $\mathcal{P}$ under tiebreaker $\tau$.
Lemma 4.2.
For every edge $(x,y)\in\mathcal{M}$ and for all tiebreakers $\tau\in\mathbb{L}$ it holds that
$state(x,y)=\textnormal{{{Fix}}}\Rightarrow(x,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau};$ (1)
$\begin{split}state(x,y)\in\{\textnormal{{BC}, {CBC}}\}&\Rightarrow\\ \exists(z,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}\colon m(z,y)\geq m(x,y);\end{split}$ (2)
$(x,y)\notin{\mathcal{M}^{\mathsf{FUN}}}\Rightarrow(x,y)\notin\mathcal{M}^{\mathsf{RV}}_{\tau}.$ (3)
Proof.
Let $\tau\in\mathbb{L}$ be an arbitrary tiebreaker. We show Equations 1 and 2 and (3) simultaneously via induction over the margin edges in order of $\tau$.
The claims are trivially true for the induction base, as $\mathcal{M}^{\mathsf{FUN}}$} is initiated as a graph with no edges. For the induction step, consider $(x,y)$ with $m(x,y)=k$. As induction hypothesis (IH) assume that Equations 1, 2 and 3 are true for all edges processed before $(x,y)$ according to $\tau$ (in particular, all edges with margin greater than $k$).
To prove Equation 1 and Equation 2, we will use the following intermediate statement, which asserts that any edge in $\mathcal{M}^{\mathsf{FUN}}$} with a final state other than CC is never rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition.
[…]
Let $(x,y)\in{\mathcal{M}^{\mathsf{FUN}}}$ with $m(x,y)=k$ be an edge with $state(x,y)\in\{\textnormal{{{Fix}}, {BC}}$, $\textnormal{{CBC}}\}$. Then $(x,y)$ cannot be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition.
Proof.
Assume towards contradiction that $state(x,y)\in\{\textnormal{{{Fix}}, {BC}}$, $\textnormal{{CBC}}\}$ but $(x,y)$ is rejected by River’s cycle condition. Then in the River diagram, there exists a path $P_{yx}\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $\operatorname{strength}(P_{yx})\geq k$, and all edges in $P_{yx}$ are processed before $(x,y)$ by $\tau$. It follows from Equation 3 of (IH) that $P_{yx}\in{\mathcal{M}^{\mathsf{FUN}}}$, as otherwise $P_{yx}\notin\mathcal{M}^{\mathsf{RV}}_{\tau}$.
Since $(x,y)\in{\mathcal{M}^{\mathsf{FUN}}}$ and $state(x,y)\in\{\textnormal{{{Fix}}},\textnormal{{BC}}$, $\textnormal{{CBC}}\}$, the edge $(x,y)$ is not updated to CC during the cycle check of any edge (line 1 of the $\mathsf{FUN}$ algorithm). Thus, there is no cycle in $\mathcal{M}^{\mathsf{FUN}}$} containing $(x,y)$ at the time it is processed. In particular, there is no path from $y$ to $x$ with strength greater than $k$. Therefore, $P_{yx}$ must contain at least one edge with margin $k$ and at least one of those margin-$k$ edges is processed after $(x,y)$ by the $\mathsf{FUN}$ algorithm. Let $e^{\prime}$ be the margin-$k$ edge which is processed last.
When $e^{\prime}$ is added to $\mathcal{M}^{\mathsf{FUN}}$}, the algorithm performs the cycle update check (line 1). Since the addition of $e^{\prime}$ closes a cycle containing $(x,y)$ in $\mathcal{M}^{\mathsf{FUN}}$} and $m(x,y)=m(e^{\prime})$, the cycle update check would update the state of $(x,y)$ to CC, contradicting the assumption. Hence, $(x,y)$ cannot be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition. ∎
We now prove the induction step for all three equations.
Equation 1: $state(x,y)=\textnormal{{Fix}}\Rightarrow(x,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$.
By Claim 4.3, the edge $(x,y)$ is not rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition. Assume towards contradiction that it is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s branching condition. Then there has to be an edge $(z,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $z\neq x$ that appears in $\tau$ before $(x,y)$. It follows by (IH) (from the contraposition of Equation 3) that $(z,y)\in{\mathcal{M}^{\mathsf{FUN}}}$. That means that $(z,y)$ is added either before or after $(x,y)$ to $\mathcal{M}^{\mathsf{FUN}}$}. If $(z,y)$ is added before $(x,y)$, then $y$ is already dominated once the algorithm processes $(x,y)$; contradicting $state(x,y)=\textnormal{{Fix}}$. Conversely, if $(z,y)$ is added after, then the state of $(x,y)$ is updated to BC in 1; contradicting $state(x,y)=\textnormal{{Fix}}$.
As a result, $(x,y)$ is not rejected by any of River’s conditions, proving $(x,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$.
Equation 2: $state(x,y)\in\{\textnormal{{BC}, {CBC}}\}\Rightarrow\exists(z,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}\colon m(z,y)\geq k$.
By Claim 4.3, the edge $(x,y)$ is not rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition. Assume $(x,y)$ is rejected by River’s branching condition. Then there has to be an edge $(z,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $m(z,y)\geq k$ and the claim holds. Otherwise, $(x,y)$ is not rejected at all by River, $(x,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ and the claim holds since $m(x,y)=k$.
Equation 3: $(x,y)\notin{\mathcal{M}^{\mathsf{FUN}}}\Rightarrow(x,y)\notin\mathcal{M}^{\mathsf{RV}}_{\tau}$
Since $(x,y)\notin{\mathcal{M}^{\mathsf{FUN}}}$, it was rejected from the $\mathsf{FUN}$ diagram – either by the branching reject check (line 1) or the cycle reject check (line 1). We show that in both cases, $(x,y)$ is also rejected from the River diagram.
First, assume $(x,y)$ was rejected by the branching reject check. Then $y$ is fixedly dominated and there exists an incoming edge to $y$ with margin $k^{\prime}>k$ in $\mathcal{M}^{\mathsf{FUN}}$}. By Equations 1 and 2 of (IH), there exists an edge $(z,y)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $m(z,y)\geq k^{\prime}$. Hence, $(x,y)$ is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s branching condition.
Now, assume $(x,y)$ was rejected by the cycle reject check. Let $U$ be the set of vertices from which there exists a path to $x$ in $\mathcal{M}^{\mathsf{FUN}}$} with strength >k>k (note that $x\in U$). Then the cycle reject check condition ensures the following.
  1. 1.
    $\mathcal{M}^{\mathsf{FUN}}$} contains a path from $y$ to $x$ with strength >k>k.
  2. 2.
    For every $u\in U$, there exists a path $P_{yu}\in{\mathcal{M}^{\mathsf{FUN}}}$ from $y$ to $u$ with $\operatorname{strength}(P_{yu})>k$ such that for every CC edge $(c,d)\in P_{yu}$, all paths $P_{dc}\in{\mathcal{M}^{\mathsf{FUN}}}$ from $d$ to $c$ with $\operatorname{strength}(P_{dc})\geq m(c,d)>k$ include vertex $y$.
We now construct a path from $y$ to $x$ in $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by iteratively extending backward from $x$, and thereby show that $(x,y)$ must be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by the cycle condition.
At each step, let $s$ be the current starting vertex of the path. If $s=y$, we are done. Otherwise, $s\in U$ and, by assumption, there exists a path from $y$ to $s$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength >k>k. Therefore, $s$ must be fixedly or cycle dominated. In both cases, we show that $s$ has an incoming edge $(c,s)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $m(c,s)>k$ such that $c\notin P$, allowing us to prepend $(c,s)$ to the path. As $y\in U$ and $U$ is finite, this backward process must eventually reach $y$, thereby completing the desired path from $y$ to $x$ in the River diagram.
Initialize $P\coloneq(x)$ as the partial path, and let $s\coloneq x$ denote the current start vertex of $P$.
$s$ is fixedly dominated.
By Equations 1 and 2 of (IH), there is some edge $(c,s)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $m(c,s)>k$. Since we only follow edges from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ in $P$ and $\mathcal{M}^{\mathsf{RV}}_{\tau}$ is a tree, $c\notin P$ as otherwise the edge $(c,s)$ would close a cycle in $\mathcal{M}^{\mathsf{RV}}_{\tau}$. We set $P\coloneq(c,s)\circ P$ and $s\coloneq c$.
$s$ is cycle dominated.
By definition of cycle dominated vertices, $s$ has at least one incoming edge in $\mathcal{M}^{\mathsf{FUN}}$} and all such incoming edges are CC. Note that $c\in U$ for all incoming edges $(c,s)\in{\mathcal{M}^{\mathsf{FUN}}}$ with $m(c,s)>k$ and, by assumption, $y$ has a path to $c$ in $\mathcal{M}^{\mathsf{FUN}}$} with strength >k>k. As a result, there exists at least one incoming edge of $s$ with margin >k>k in $\mathcal{M}^{\mathsf{FUN}}$}, and all such incoming edges are on a path from $y$ with strength >k>k.
If one such edge $(c,s)$ is also in the River diagram, then $c\notin P$ (as otherwise the edge $(c,s)$ would close a cycle in $\mathcal{M}^{\mathsf{RV}}_{\tau}$) and we set $P\coloneq(c,s)\circ P$ and $s\coloneq c$.
Otherwise, every such edge $(c,s)\in{\mathcal{M}^{\mathsf{FUN}}}$ with margin >k>k is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$.
  • If $(c,s)$ is rejected by River’s branching condition, then in the River diagram, there is $(c^{\prime},s)\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ with $m(c^{\prime},s)\geq m(c,s)>k$. By the contraposition of Equation 3 of (IH), $(c^{\prime},s)\in{\mathcal{M}^{\mathsf{FUN}}}$ and we set $P\coloneq(c^{\prime},s)\circ P$ and $s\coloneq c^{\prime}$.
  • If every such edge $(c,s)$ is rejected by River’s cycle condition, then in the River diagram, there is a path $P_{sc}\in\mathcal{M}^{\mathsf{RV}}_{\tau}$ from $s$ to $c$ with strength $\geq m(c,s)>k$. By the contraposition of Equation 3 of (IH), $P_{sc}\in{\mathcal{M}^{\mathsf{FUN}}}$. Now, since $(c,s)$ is a CC edge on a path from $y$ to $s$ and $\operatorname{strength}(P_{sc})\geq m(c,s)>k$, our assumption implies $y\in P_{sc}$.
    In conclusion, there exists a path $P_{sc}$ in the River diagram which includes $y$. In this case, $y$ already has an incoming edge in the River diagram $\mathcal{M}^{\mathsf{RV}}_{\tau}$ and $(x,y)$ is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by the branching condition.
Therefore, either $(x,y)$ is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by Rivers’ branching condition, or there exists a path from $y$ to $x$ in $\mathcal{M}^{\mathsf{RV}}_{\tau}$ with strength >m(x,y)>m(x,y), because of which $(x,y)$ is rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau}$ by River’s cycle condition.
This proves the induction step for Equations 1, 2 and 3, and therefore shows the claim. ∎ With this, we can now prove the forward directionof the correctness theorem.
Theorem 4.4 (Forward Direction of Main Theorem).
For every preference profile $\mathcal{P}$ and every alternative $a\in A$:
$a\in\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})$ $\Rightarrow a\text{ is cycle or not dominated in }{\mathcal{M}^{\mathsf{FUN}}}(\mathcal{P}).$
Proof.
We show the contraposition of the claim. For that, recall that by definition every alternative that is neither cycle dominated nor not dominated, must be fixedly dominated. The contraposition can therefore be formulated as $a\notin\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})\Leftarrow a$ is fixedly dominated. Suppose $a$ is fixedly dominated. Then, by definition, it has either an incoming edge labeled Fix, or at least one incoming edge labeled BC or CBC in $\mathcal{M}^{\mathsf{FUN}}$}. By Lemma 4.2, there exists some incoming edge to $y$ in every universe. Therefore, $a\notin\mathsf{RV}(\mathcal{P},\tau)$ and $a\notin\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})$. ∎
4.2 Backward Direction: Cycle Edges and a Winning Certificate
In this section, we show that every alternative $a$ identified as an $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winner by the $\mathsf{FUN}$ algorithm is indeed a River winner in at least one tiebreaker universe. To that end, we present a procedure that extracts a River diagram from $\mathcal{M}^{\mathsf{FUN}}$} that certifies $a$ as a winner. We also describe how to construct the tiebreaking order that defines this universe.
Deriving a Potential Certificate for an Alternative
Recall that, for a given preference profile $\mathcal{P}$ and tiebreaker $\tau$, the River Method computes a tree rooted in some alternative $a$: the River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$. This tree certifies $a$ as the winner by containing, for every other alternative $b\neq a$, a path from $a$ to $b$.
In a similar fashion, we now describe, given the Fused-Universe diagram $\mathcal{M}^{\mathsf{FUN}}$} and an alternative $a$, how to compute a tree $T({\mathcal{M}^{\mathsf{FUN}}},a)$ rooted at $a$. If $a$ is an $\operatorname{\mathsf{RV\mbox{-}PUT}}$ winner, this tree will serve as a certificate: it coincides with the River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau)$ for some tiebreaker $\tau$ under which $a$ is the unique River winner. If $\mathcal{M}^{\mathsf{FUN}}$} is clear from the context, we denote the tree $T({\mathcal{M}^{\mathsf{FUN}}},a)$ with $T^{a}$.
The algorithm computing $T^{a}$ is a variant of Prim’s algorithm for computing a Minimum Spanning Tree (MST) (Prim 1957; Jarník 1930; Dijkstra 1959), therefore named DirectedMaxPrim:
  1. 1.
    Initialize $T=(\{a\},\emptyset)$.
  2. 2.
    As long as there are crossing edges $E^{c}=\{\,(u,v)\in{\mathcal{M}^{\mathsf{FUN}}}\mid u\in T\land v\notin T\,\}$, add to $T$ an edge from $E^{c}$ with maximum margin.
  3. 3.
    Return $T^{a}=T$.
An efficient implementation uses a priority queue such as Fibonacci Heaps (Fredman and Tarjan 1987) to select the edges in step 2. For a $\mathsf{FUN}$ diagram with $|A|$ alternatives and $|E|$ edges, DirectedMaxPrim has a runtime in $O(|E|+|A|\log|A|)$ (Cormen et al. 2022).
Note that since the algorithm is initialized at $a$ and only adds edges from vertices in $T$ to vertices outside $T$, the resulting $T^{a}$ is a (not necessarily spanning) tree rooted in $a$. We now state two properties of $T^{a}$. The first establishes the existence of strong paths within the tree.
Lemma 4.5 ($\star$).
If a $\mathsf{FUN}$ diagram $\mathcal{M}^{\mathsf{FUN}}$} contains a path $P_{ab}$ from $a$ to $b$, then $T^{a}$ contains a path $P_{ab}^{\prime}$ from $a$ to $b$ with $\operatorname{strength}(P_{ab}^{\prime})\geq\operatorname{strength}(P_{ab})$.
Proof.
Fix ${\mathcal{M}^{\mathsf{FUN}}},a,b,P_{ab}$ and let $T^{a}$ be the output of DirectedMaxPrim. Assume towards contradiction that $T^{a}$ contains no such path $P_{ab}^{\prime}$ and distinguish two cases:
  1. 1.
    There is no path from $a$ to $b$ in $T^{a}$. Let $u$ be the last vertex on $P_{ab}$ that is reachable from $a$ in $T^{a}$ and let $v$ be the successor of $u$ on $P_{ab}$. But then $(u,v)$ is a crossing edge and thus DirectedMaxPrim cannot have terminated with $T^{a}$ as output due to the loop condition in step 2.
  2. 2.
    The unique path $P_{ab}^{\prime}$ from $a$ to $b$ in $T^{a}$ has $\operatorname{strength}(P_{ab}^{\prime})<\operatorname{strength}(P_{ab})$. Let $e^{\prime}$ be an edge of minimum margin in $P_{ab}^{\prime}$. Recall how DirectedMaxPrim iteratively builds $T^{a}$. Inspect the moment, in which the algorithm picks $e^{\prime}$ as edge with maximum margin from the current set of crossing edges $E^{c}$ and adds it to $T$. At this moment, $b$ is not yet reachable from $a$ in $T$. Thus there is a first edge $e\in E^{c}$ on $P_{ab}$ that is not in $T$. However, from $\operatorname{strength}(P_{ab}^{\prime})<\operatorname{strength}(P_{ab})$ we can infer $\operatorname{strength}(e^{\prime})<\operatorname{strength}(e)$, which contradicts that $e^{\prime}$ has maximum margin in $E^{c}$.
Figure 2: Exemplary illustration for the proof of Lemma 4.6. Dotted arrows represent paths, and the shaded region indicates the current tree $T$. All solid arrows are edges in the Fused-Universe diagram $\mathcal{M}^{\mathsf{FUN}}$}. From left to right: (i) before DirectedMaxPrim picks $(d,s)$ from the current set of crossing edges, with $d^{L}\notin T$; (ii) after $(d,s)$ is added to $T$ and a path $P_{sd^{L}}$ exists in $\mathcal{M}^{\mathsf{FUN}}$}; (iii) when DirectedMaxPrim explores $d^{L}$ via a path from $s$.
The next lemma explains how an edge can belong to the certificate tree even when a higher-margin edge to the same alternative exists. See Figure 2 for an illustration. Specifically, if an edge $(d,s)\in{\mathcal{M}^{\mathsf{FUN}}}$ is chosen for the tree $T^{a}$ despite the existence of a higher-margin edge $(d^{L},s)\in{\mathcal{M}^{\mathsf{FUN}}}$, then $(d^{L},s)$ must be an CC edge. This implies the existence of a strong enough path from $s$ to $d^{L}$ in $\mathcal{M}^{\mathsf{FUN}}$}, which can be used in place of the direct edge $(d^{L},s)$.
Lemma 4.6 ($\star$).
Let $\mathcal{M}^{\mathsf{FUN}}$} be a $\mathsf{FUN}$ diagram with margins $m$. Let $a$ be an alternative and let $s,d,d^{L}$ be different alternatives reachable from $a$ in $\mathcal{M}^{\mathsf{FUN}}$}. If $\mathcal{M}^{\mathsf{FUN}}$} contains edges $(d,s),(d^{L},s)$ with $m(d,s)<m(d^{L},s)$ and $T^{a}$ contains $(d,s)$, then $T^{a}$ contains a path $P_{sd^{L}}$ from $s$ to $d^{L}$ with $\operatorname{strength}(P_{sd^{L}})\geq m(d^{L},s)$.
Proof.
Fix ${\mathcal{M}^{\mathsf{FUN}}},a,s,d,d^{L}$ with $(d,s),(d^{L},s)\in{\mathcal{M}^{\mathsf{FUN}}}$, $m(d,s)<m(d^{L},s)$ and $(d,s)\in T^{a}$.
Inspect the moment, in which DirectedMaxPrim picks $(d,s)$ as edge with maximum margin from the current set of crossing edges $E^{c}$ and adds it to $T$. At that moment, $d^{L}\notin T$; otherwise $(d^{L},s)$ would be a crossing edge with higher margin. For an illustration, see Figure 2 (i).
Since $(d,s),(d^{L},s)\in{\mathcal{M}^{\mathsf{FUN}}}$, it follows that $(d^{L},s)$ must be a CC edge: otherwise $s$ would be fixedly dominated by an edge with margin at least $m(d^{L},s)$ and no edge with margin lower than $m(d^{L},s)$ would have been added to $\mathcal{M}^{\mathsf{FUN}}$}. By definition of CC edges there exists a path $P_{sd^{L}}^{\prime}$ from $s$ to $d^{L}$ in $\mathcal{M}^{\mathsf{FUN}}$} with $\operatorname{strength}(P_{sd^{L}}^{\prime})\geq m(d^{L},s)$. (See Figure 2 (ii).)
From $m(d^{L},s)>m(d,s)$ follows that all edges in $P_{sd^{L}}^{\prime}$ have margin greater than that of any other edge in $E^{c}$. Therefore, DirectedMaxPrim explores $d^{L}$ via a path $P_{sd^{L}}$ starting in $s$ with $\operatorname{strength}(P_{sd^{L}})\geq\operatorname{strength}(P_{sd^{L}}^{\prime})$ before considering any other edges in $E^{c}$. This is illustrated in Figure 2 (iii). ∎
From the certificate tree $T^{a}$, we derive a tiebreaker $\tau^{a}$ that prioritizes the edges in $T^{a}$ over any other edges with the same margin that are not in $T^{a}$. We will show in Theorem 4.9 that whenever $a$ is not dominated or cycle dominated in $\mathcal{M}^{\mathsf{FUN}}$}, running River with this tiebreaker will result in a River diagram $\mathcal{M}^{\mathsf{RV}}(\mathcal{P},\tau^{a})$ which exactly matches $T^{a}$.
Definition 4.7.
Let $\mathcal{P}$ be a preference profile and $\mathcal{M}(\mathcal{P})$ the corresponding margin graph. A certificate tiebreaker $\tau^{a}$ for an alternative $a$ with certificate tree $T^{a}$ is an ordering of all edges in $\mathcal{M}(\mathcal{P})$ that fulfills the following two properties:
  1. 1.
    Edges are ordered by descending margin.
  2. 2.
    Among edges with identical margin, those that are in $T^{a}$ come before those not in $T^{a}$.
Confirming $\operatorname{\mathsf{RV\mbox{-}PUT}}$-Winners Using the Certificate Tree
Note that $T^{a}$ is not necessarily a spanning tree, as some alternatives may not be reachable from $a$ in $\mathcal{M}^{\mathsf{FUN}}$}. However, if $a$ is cycle dominated or not dominated, then the certificate tree $T^{a}$ is a spanning tree over all alternatives:
Lemma 4.8 ($\star$).
If an alternative $a$ is cycle dominated or not dominated in $\mathcal{M}^{\mathsf{FUN}}$} then the certificate tree $T^{a}$ is a spanning tree of $\mathcal{M}$ rooted in $a$.
Proof.
For the definitions of cycle and not dominated, refer to Table 1. We distinguish between $a$ being not dominated and $a$ being cycle dominated, and show that $a$ can reach every other alternative in $\mathcal{M}^{\mathsf{FUN}}$} in both cases.
Case 1: $a$ is not dominated in $\mathcal{M}^{\mathsf{FUN}}$}.
This means $a$ has no incoming edge in $\mathcal{M}^{\mathsf{FUN}}$}. Consider any vertex $x\in A\setminus\{a\}$. We walk backward from $x$ using only incoming edges labeled Fix, BC, or CBC until one of the following happens:
  1. (i)
    We reach $a$: then clearly $a$ can reach $x$ via the reverse of this walk.
  2. (ii)
    We visit a vertex twice: then we have found a cycle in $\mathcal{M}^{\mathsf{FUN}}$} composed solely of Fix, BC, or CBC edges. But this contradicts the $\mathsf{FUN}$ algorithm’s handling of cycles: any such cycle must contain (i) a unique weakest edge, which would have been rejected by the cycle reject check; or (ii) multiple weakest edges, all of which would have been updated to CC in the cycle update check.
  3. (iii)
    We get stuck at a vertex $x^{\prime}$ where all incoming edges in $\mathcal{M}^{\mathsf{FUN}}$} are labeled CC. Then $x^{\prime}$ is cycle dominated, and we distinguish two subcases:
    • If $(a,x^{\prime})\in\mathcal{M}$, then $(a,x^{\prime})$ was not rejected by the branching condition (since $x^{\prime}$ is not fixedly dominated), and not rejected by the cycle condition (since $a$ has no incoming edge in $\mathcal{M}^{\mathsf{FUN}}$}). Hence, $(a,x^{\prime})\in{\mathcal{M}^{\mathsf{FUN}}}$.
    • If $(x^{\prime},a)\in\mathcal{M}$, then since $a$ is not dominated, $(x^{\prime},a)\notin{\mathcal{M}^{\mathsf{FUN}}}$ and it was rejected by the cycle reject check. That implies that there exists a path from $a$ to $x^{\prime}$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength at least $m(x^{\prime},a)$.
    In either subcase, $a$ can reach $x^{\prime}$ and thereby $x$.
Case 2: $a$ is cycle dominated in $\mathcal{M}^{\mathsf{FUN}}$}.
Then $a$ has at least one incoming edge in $\mathcal{M}^{\mathsf{FUN}}$} and all such edges are CC. Consider any vertex $x\in A\setminus\{a\}$. As in Case 1, we walk backward from $x$ using only incoming Fix, BC, or CBC edges until one of the following happens:
  1. (i)
    We reach $a$: then clearly $a$ can reach $x$ via the reverse of this walk.
  2. (ii)
    We visit a vertex twice: then we have found a cycle in $\mathcal{M}^{\mathsf{FUN}}$} composed solely of Fix, BC, or CBC edges. As in Case 1, this contradicts the $\mathsf{FUN}$ algorithm’s handling of cycles (see Case 1 (ii)).
  3. (iii)
    We get stuck at a vertex $x^{\prime}$ where all incoming edges in $\mathcal{M}^{\mathsf{FUN}}$} are labeled CC. Then $x^{\prime}$ is cycle dominated, and we distinguish two subcases:
    $(a,x^{\prime})\in\mathcal{M}$.
    If $(a,x^{\prime})\in{\mathcal{M}^{\mathsf{FUN}}}$, then $a$ can reach $x^{\prime}$.
    Otherwise, $(a,x^{\prime})\notin{\mathcal{M}^{\mathsf{FUN}}}$. Since $x^{\prime}$ is not fixedly dominated, $(a,x^{\prime})$ was not rejected from $\mathcal{M}^{\mathsf{FUN}}$} by the branching reject check, and thus was rejected by the cycle reject check. This implies a path from $x^{\prime}$ to $a$ in $\mathcal{M}^{\mathsf{FUN}}$} of strength $m(a,x^{\prime})$, and for every CC edge $(c,d)$ on such a path, the cycle reject condition ensures that all paths from $d$ to $c$ with margin $\geq m(c,d)$ go through $x^{\prime}$. In particular, the edge $(c,a)$ at the end of such a path must satisfy this condition: since $(c,a)$ is CC (as $a$ is cycle dominated), there exists a path from $a$ to $c$ with strength $\geq m(c,a)$ and by the condition above, this path must pass through $x^{\prime}$. Hence, $a$ can reach $x^{\prime}$.
    $(x^{\prime},a)\in\mathcal{M}$.
    If $(x^{\prime},a)\in{\mathcal{M}^{\mathsf{FUN}}}$, then it must be CC since $a$ is cycle dominated. This implies that there exists a path from $a$ to $x^{\prime}$ in $\mathcal{M}^{\mathsf{FUN}}$}.
    Otherwise, $(x^{\prime},a)\notin{\mathcal{M}^{\mathsf{FUN}}}$. Since $a$ is not fixedly dominated, $(x^{\prime},a)$ was not rejected from $\mathcal{M}^{\mathsf{FUN}}$} by the branching reject check, and thus was rejected by the cycle reject check. This implies that there is a path from $a$ to $x^{\prime}$ in $\mathcal{M}^{\mathsf{FUN}}$}.
    In either subcase, $a$ can reach $x^{\prime}$ and thereby $x$.
In both Case 1 and Case 2, $a$ has a directed path in $\mathcal{M}^{\mathsf{FUN}}$} to every other alternative. Lemma 4.5 implies that $T^{a}$ is a spanning tree of $\mathcal{M}^{\mathsf{FUN}}$} rooted in $a$. ∎ With this, we are ready to prove the backward direction of Theorem 4.1.
Theorem 4.9Backward Direction of Main Theorem. (Backward Direction of Main Theorem).
For every preference profile $\mathcal{P}$ and every alternative $a\in A$:
$a\in\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})$ $\Leftarrow a\text{ is cycle or not dominated in }{\mathcal{M}^{\mathsf{FUN}}}(\mathcal{P}).$
Proof.
Let $\mathcal{P}$ be a preference profile and let $a\in A$ be an alternative. Recall the definition of $\operatorname{\mathsf{RV\mbox{-}PUT}}$ which states $a\in\operatorname{\mathsf{RV\mbox{-}PUT}}(\mathcal{P})\Leftrightarrow\exists\tau\in\mathbb{L}\colon\mathsf{RV}(\mathcal{P},\tau)=\{a\}$. Thus, we have to show:
$\exists{\tau}\in\mathbb{L}:\mathsf{RV}(\mathcal{P},\tau)=\{a\}\Leftarrow$ $a\text{ is cycle dominated or}$
$\text{ not dominated in }{\mathcal{M}^{\mathsf{FUN}}}(\mathcal{P}).$
Assume $a$ is either cycle dominated or not dominated in $\mathcal{M}^{\mathsf{FUN}}$}. Then by Lemma 4.8, the certificate tree $T^{a}$ of $a$ is a spanning tree. Let $\tau^{a}$ be the corresponding certificate tiebreaker. We prove that $\mathsf{RV}(\mathcal{P},\tau^{a})=\{a\}$ by establishing the stronger result that the River diagram under $\tau^{a}$ is exactly the certificate tree rooted in $a$, i. e., $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}=T^{a}$.
Since $T^{a}$ and $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ are trees on the same vertex set, it suffices to show that one edge set is subset of the other. We show $T^{a}\subseteq\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ by induction over the edges of $\mathcal{M}^{\mathsf{FUN}}$} in the order $\tau^{a}=(e_{0},e_{1},\ldots,e_{\ell})$. For the induction base consider the edge $e_{0}$ and assume $e_{0}\in T^{a}$. Since $e_{0}$ is the first edge processed by River, $e_{0}$ cannot be rejected by $(Br)$ or $(Cy)$, and consequently $e_{0}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$.
Now, let $0<i<\ell$ and assume as induction hypothesis (IH) for $0\leq j<i$ that $e_{j}\in T^{a}$ implies $e_{j}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$. We use this to show that $e_{i}\in T^{a}$ implies $e_{i}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$. Denote the endpoints of the edge as $(x,y)=e_{i}$ and assume towards contradiction that $(x,y)\in T^{a}$ but $(x,y)\notin\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$. This implies that $(x,y)$ was rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ by River’s branching condition or cycle condition. We distinguish accordingly:
Case 1: $(x,y)$ is rejected by the branching condition.
Consequently, there exists an edge $(x^{L},y)\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ with $m(x,y)\leq m(x^{L},y)$. Since $T^{a}$ is a tree with at most one incoming edge per vertex and $(x,y)\in T^{a}$, we conclude $(x^{L},y)\notin T^{a}$.
If $m(x^{L},y)=m(x,y)$, then by definition of $\tau^{a}$, $(x,y)$ is processed by River before $(x^{L},y)$. But then $(x,y)$ cannot be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ by $(Br)$ because of $(x^{L},y)$, since $(x^{L},y)$ has not even been processed yet.
Therefore, we can assume $m(x,y)<m(x^{L},y)$. Now, observe that $(x,y)\in{\mathcal{M}^{\mathsf{FUN}}}$ and $(x^{L},y)\in{\mathcal{M}^{\mathsf{FUN}}}$. It follows from Lemma 4.6 with $d=x$, $d^{L}=x^{L}$, and $s=y$ that $T^{a}$ includes a path from $y$ to $x^{L}$ with strength larger $m(x,y)$. By (IH) $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ includes the same path and thus cannot include $(x^{L},y)$ – a contradiction.
Case 2: $(x,y)$ is rejected by the cycle condition.
Consequently, there is a path $P_{yx}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ with $m(x,y)\leq\operatorname{strength}(P_{yx})$. Since $T^{a}$ is acyclic, $P_{yx}$ cannot be fully contained in $T^{a}$, i. e., there must be at least one edge $(u,v)\in P_{yx}$ with $(u,v)\notin T^{a}$ and $m(x,y)\leq m(u,v)$. We show that this implies $(u,v)\notin\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$, which contradicts $P_{yx}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$.
If $m(x,y)=m(u,v)$, then by definition of $\tau^{a}$, $(x,y)$ is processed by River before $(u,v)$. But then $(x,y)$ cannot be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ by $(Cy)$ because of $P_{yx}$, since $(u,v)$ has not even been processed yet and consequently $P_{yx}$ is not yet in $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$.
Therefore, we can assume $m(x,y)<m(u,v)$. Now, from $(u,v)\notin T^{a}$ follows that either (1) there is some other edge $(u^{\prime},v)\in T^{a}$ which is chosen instead of $(u,v)$, or (2) there is no edge towards $v$ in $T^{a}$. We distinguish accordingly:
Case 2.1
Assume there is some other edge $(u^{\prime},v)\in T^{a}$ which is chosen instead of $(u,v)$. Observe that $(u^{\prime},v)\in{\mathcal{M}^{\mathsf{FUN}}}$ and $(u,v)\in{\mathcal{M}^{\mathsf{FUN}}}$. If $m(x,y)<m(u^{\prime},v)$, then $(u^{\prime},v)\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ follows from (IH). In that case, $(u,v)$ would be rejected by $(Br)$ from $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$, i. e., $(u,v)\notin\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$. If $m(u^{\prime},v)\leq m(x,y)$, then $m(u^{\prime},v)<m(u,v)$ and it follows from Lemma 4.6 with $d=u^{\prime}$, $d^{L}=u$, and $s=v$ that $T^{a}$ includes a path from $v$ to $u$ with strength larger $m(u,v)>m(x,y)$. By (IH), $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ includes the same path and thus cannot include $(u,v)$.
Case 2.2
Assume there is no edge towards $v$ in $T^{a}$. In that case, $v$ must be the root $a$, and we have to show $(u,a)\notin\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$. As $a$ has only incoming CC edges in $\mathcal{M}^{\mathsf{FUN}}$}, $(u,a)$ has to be a CC edge. By the definition of CC there exists a path $P_{au}$ from $a$ to $u$ in $\mathcal{M}^{\mathsf{FUN}}$} with $\operatorname{strength}(P_{au})\geq m(u,a)$. It follows from Lemma 4.5 that $T^{a}$ contains a path $P_{au}^{\prime}$ from $a$ to $u$ with $\operatorname{strength}(P_{au}^{\prime})\geq\operatorname{strength}(P_{au})$
As $\operatorname{strength}(P_{au})>m(x,y)$ (IH) implies that $P_{au}^{\prime}\in\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ and $(u,a)$ would be rejected from $\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$ by $(Cy)$, i. e., $(u,v)=(u,a)\notin\mathcal{M}^{\mathsf{RV}}_{\tau^{a}}$.
5 Experiments
Figure 3: A scatterplot of all recorded running times of the $\mathsf{FUN}$ algorithm versus $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$ as implemented in pref-voting by (Mattei and Walsh 2013) over synthetic election with different numbers of alternatives and voters on a logarithmic scale. Note that after 3 timeouts, no more running times where recorded for that $m$.
Figure 4: Average running time of the $\mathsf{FUN}$ algorithm versus Stable Voting, Beat Path, and Split Cycle as implemented in pref-voting by (Mattei and Walsh 2013) over single runs per 20 synthetic elections with different numbers of alternatives and voters on a logarithmic scale.
To assess the practical performance of the $\mathsf{FUN}$ algorithm, we conducted preliminary experiments comparing its running time to that of other voting methods. In particular, we evaluated River $\mathsf{FUN}$ against River with brute-force PUT ($\operatorname{\mathsf{RV\mbox{-}PUT}}$), Ranked Pairs with brute-force PUT ($\operatorname{\mathsf{RP\mbox{-}PUT}}$), as well as the polynomial-time methods Split Cycle, Stable Voting, and Beat Path, using implementations from the pref_voting Python library (Holliday and Pacuit 2025).
5.1 Setup
We followed the recommendations by (Szufa et al. 2025) and generated synthetic elections using the normalized Mallows model with dispersion parameter $\text{norm-}\phi=0.7$, which simulates elections that resemble real-world data. We focused only on elections without a Condorcet winner, since all considered methods are Condorcet consistent and first check for Condorcet winners, therefore taking the same amount of time to return the winner. We generated elections for all combinations of:
  • number of alternatives $m\in\{5,6,\dots,50\}$, and
  • number of voters $n\in\{10,50,100,200\}$.
For each $(m,n)$ pair, we generated elections until we obtained 20 elections without a Condorcet winner or until we discarded 50,000 samples. For the parameter combinations $m=5$ with $n=50$, for $m\in\{5,6,7,8,9\}$ with $n=100$, as well as for $m\in\{5,\dots,15\}$ with $n=200$ where we could not generate 20 suitable elections, we excluded them from the results.
All voting methods were run sequentially in randomized order across all eligible elections. To keep computing time within a feasible timeframe, we applied the following timeouts:
  • 5 minutes for polynomial-time methods (Split Cycle, Stable Voting, Beat Path and River $\mathsf{FUN}$), and
  • 30 minutes for $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$.
If $\operatorname{\mathsf{RV\mbox{-}PUT}}$ or $\operatorname{\mathsf{RP\mbox{-}PUT}}$ exceeded the timeout $3$ times for one $m$ they where not run again for that $m$. Since pilot experiments indicated that both methods would exceed the timeout at small election sizes, we limited their experiments to $m\in\{5,6,\dots,12\}$. The polynomial-time methods where not rerun for $(m,n)$ once if they had exceeded their timeout.
The experiments were run on a $2.30$ GHz core with 64GB RAM under Ubuntu 24.04.2 LTS, using CPython 3.12.3 with pref-voting 1.16.21, multiprocess 0.70.18, and networkx 3.5.
5.2 Goals
The goal of our experiments was twofold:
Scalability
Determine whether the $\mathsf{FUN}$ algorithm makes computing $\operatorname{\mathsf{RV\mbox{-}PUT}}$ feasible on instances where brute-force methods fail.
Performance
Compare $\mathsf{FUN}$ ’s running time to that of other common polynomial-time voting methods.
5.3 Results
Our findings are illustrated in Figure 3 and Figure 4.
River $\mathsf{FUN}$ vs $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$
Our results confirm that the $\mathsf{FUN}$ algorithm makes River with Parallel-Universe Tiebreaking computationally feasible even for moderately sized elections. As shown in Figure 3, both $\operatorname{\mathsf{RV\mbox{-}PUT}}$ and $\operatorname{\mathsf{RP\mbox{-}PUT}}$ exceeded the 30-minute timeout on more than three instances for $m\geq 7$. In contrast, the $\mathsf{FUN}$ algorithm consistently computed the River outcome in under 0.1 seconds—on the same instances.
River $\mathsf{FUN}$ vs. Other Polynomial-Time Methods
$\mathsf{FUN}$ ’s performance is comparable to that of Split Cycle, Beat Path, and Stable Voting (Figure 4). The implementations of Split Cycle and Beat Path in (Mattei and Walsh 2013) both use the Floyd–Warshall algorithm, which explains their nearly identical performance. While $\mathsf{FUN}$ is slightly slower than the other methods for $m\leq 16$, its performance surpasses Stable Voting from $m\geq 17$ and surpasses/ is on par with Beat Path and Split Cycle from $m\geq 20$. Note that for $n=10$, its running time closely matches that of Beat Path and Split Cycle.
We also observed that $\mathsf{FUN}$ ’s performance varies with the number of voters: when $n$ is small relative to $m$, the computation time increases. We conjecture this is due to a higher number of nearly tied alternatives, causing more edges to be retained in the $\mathsf{FUN}$ diagram.
Finally, we observed that Stable Voting exhibits greater variance in its running time compared to the other methods, which show smooth scaling. This is likely due to Stable Voting’s recursive definition, which can require many repetitions of winner computation depending on the profile structure.
6 Conclusion
We presented a polynomial-time algorithm for computing the PUT (Parallel-Universe Tiebreaking) variant of River. Our $\mathsf{FUN}$ algorithm exploits River’s branching condition to maintain a compact diagram capturing all possible River outcomes across tiebreakers. This highlights the computational benefit of River’s simple decision process, especially in contrast to Ranked Pairs, whose PUT variant is $\mathtt{NP}$-hard.
While the $\mathsf{FUN}$ diagram compactly encodes all possible River outcomes, it may include edges that do not appear in any River diagram Our algorithm only removes edges which are rejected by River in all universes for the same reason. Thus it does not remove edges rejected sometimes by branching and sometimes by cycle condition. While our algorithm is sufficient for correctness, further work could refine this to compute the exact union of all River diagrams.
As $\operatorname{\mathsf{RV\mbox{-}PUT}}$ defines a new social choice function, an axiomatic study could be of interest. In particular in comparison to PUT-Ranked Pairs and River in terms of winning set containment and axioms such as clone independence.
Preliminary experiments indicate that $\mathsf{FUN}$ scales to instances where brute-force PUT fails and is competitive with well-regarded polynomial-time rules. A broader empirical study on synthetic and real data, including comparison to the optimized $\operatorname{\mathsf{RP\mbox{-}PUT}}$ implementation by Wang et al. [^1] We thank an anonymous AAAI 2026 reviewer for this pointer., is promising future work.
Acknowledgments
The project on which this report is based was funded by the Federal Ministry of Research, Technology and Space under the funding code “KI-Servicezentrum Berlin-Brandenburg” 16IS22092. Responsibility for the content of this publication remains with the author (Michelle Döring).